# DepthVanish: Optimizing Adversarial Interval Structures for Stereo-Depth-Invisible Patches

### v1

@@ 6.5 Ablation Study
Figure 9: Attack performance of RAFT-Stereo under different $\alpha$ and $\beta$. Evaluated on KITTI dataset.
In this section, we conduct ablation analysis on the DepthVanish attack to assess the impact of the hyperparameters $\alpha$ and $\beta$ in the objective function (Eq. (9)). As shown in Fig. 9, both parameters are critical for optimal attack performance. Specifically, it can be seen that the performance degraded significantly when $\alpha=0$, i.e., the $\mathcal{L}_{entropy}$ is removed from Eq. (9), which highlights the importance of the clear interval for attack effectiveness. Moreover, the total variation constraint $\mathcal{L}_{tv}$ is also important where a clear performance degradation can be observed when $\beta$ decreases below 9. In summary, the collaboration of entropy and total variation regularizations effectively enables our DepthVanish patch to achieve the superior attack performance.
@@ 7 Conclusion
In this work, we present DepthVanish, a significant advancement of an adversarial patch attack that jointly optimizes texture and interval structure to target stereo depth systems. By thoroughly analyzing the effect of interval spacing over repeated texture elements, we provide a novel insight for boosting the attack effectiveness on traditional repeated texture. Unlike previous approaches limited to digital environments, our method effectively “disappears” objects in both digital and physical settings when tested against state-of-the-art models and commercial RGB-D cameras. These findings expose critical vulnerabilities in current depth estimation technologies, highlighting the urgent need for robust defenses in safety-critical autonomous systems.
@@ Appendix A Experimental Environment
All of the experiments are conducted on a server with AMD EPYC 9554 64-core Processor and an NVIDIA L40 GPU, running Ubuntu 22.04. During the physical evaluation, the patch is first printed out with an EPSON L18050 printer, e.g., a patch of physical size $(u,v)=(0.891m,1.26m)$ is printed out by filling A3 size paper.
@@ B.1 Calculation for Digital Deployment
During the digital experiments, the patches are first digitally deployed into the scene. As our aim is to achieve physical attack, it requires the digital deployment follows the physical constraints. Fortunately, such physical-constraint digital patch deployment can be realized with the calibration information provided by the dataset. We show the detail calculation for KITTI dataset in the follow.
For KITTI dataset, we suppose the patch’s physical size and depth are predefined with regard to the camera 0 (the reference camera). Given the patch with physical size of $(w_{p},h_{p})$, we intend to deploy it into a KITTI scene with physical depth $e$. 1. Get the corresponding calibration information for the scene from ‘calib_cam_to_cam’ folder provided by KITTI scene flow dataset. 2. Retrieve the three rectified calibration matrix $P\_rect\_02$, $P\_rect\_03$, $R\_rect\_00$. 3. Specify the physical shifting $(x_{shift},y_{shift})$ (in meters) of the patch center with regard to the camera 0 principal axis. 4. Set the homogeneous coordinates for the corners of the patch as: $top\_left=(-w_{p}/2+x_{shift},-h_{p}/2+y_{shift},e,1)$, $top\_right=(w_{p}/2+x_{shift},-h_{p}/2+y_{shift},e,1)$, $bottom\_left=(-w_{p}/2+x_{shift},h_{p}/2+y_{shift},e,1)$, $bottom\_right=(-w_{p}/2+x_{shift},h_{p}/2+y_{shift},e,1)$. 5. For the pixel coordinates of the patch’s corner in the right stereo image, get them with $(P\_rect\_02\times R\_rect\_00)\cdot c$ where $c$ is the corners homogeneous coordinates. 6. Similarly, get the corresponding pixel coordinates for the left stereo image with $(P\_rect\_03\times R\_rect\_00)\cdot c$. 7. Finally, the patch is deployed into the scene by applying perspective transformation to fit the patch into the calculated pixel region.
For the calculation of DrivingStereo, the process is the same as KITTI except for the reference camera is camera 1 thus $P\_rect\_101$, $P\_rect\_103$ and $R\_rect\_101$ from their calibration file is adopted.
@@ B.2 Physical Stereo Image Capture
During the experiments of physical evaluation of our patch, we have to manually capture the scene due to the lack of RGB stereo capturing device. We adopt the i3DStreoid mobile application which is specifically designed to facilitate the capturing of stereo images with mobile phone of model iPhone 14 pro. In specific, a cutting board of A3 size is utilized to place the mobile phone for a $30cm$ baseline simulation. Then at each place a picture is taken by the mobile phone as one of the stereo image. However, we do aware the inaccuracy of such stereo image capturing process that is also we further test our patch with commercial RGB-D camera, i.e., IntelRealSense D435i.
@@ B.3 Setting for Attack Targets
For the PSMNet, DeepPruner and AANet, we follow the setup of Stereoscopic [1] and directly adopt the API provided at https://github.com/alexklwong/stereoscopic-universal-perturbations.git. As for the RAFT-Stereo and STTR, we use the official code release and integrate them following Stereoscopic code. And the checkpoint pretrianed on KITTI dataset for both models, i.e., ‘raftstereo-sceneflow.pth’ and ‘kitti_finetuned_model.pth.tar’, are adopt for experiments. Note that we keep setting $k=3$ for D1-error metric calculation during the evaluation for all the attack taregt models.
@@ C.1 Optimize under Different Setup
We conduct experiments to test the influence of the patch physical size and depth. As shown in Fig. 2, we test physical depth ranges from $5m$ to $21m$ and different physical size by setting the patch as a scaled size of $(0.891m,1.26m)$. In general, we see the patch keeps be attack effective across different physical size and depth. Empirically, a patch will show great disturbance when the D1-error larger than 80 which physically represents an disappeared object for the stereo system.
@@ C.2 Attack Transferability Results
As we have show our DepthVanish patch enjoys strong attack transferability over different stereo estimation models, we further visualize such merit in Fig. 3 and Fig. 4 on KITTI and DrivingStereo dataset respectively. Although it can be seen that there are some unsatisfying results on KITTI dataset, we can observe our DepthValish patch keeps successfully attacking all the models with different transferability.
@@ Appendix D Limitation
Our research has several important limitations regarding the attack methodology. Due to limited testing scope, we cannot guarantee the robustness of our attack methods against the full spectrum of existing DNN architectures for stereo depth estimation. Additionally, we have not evaluated our attacks against existing robustified methods or defense mechanisms that may be implemented in real-world systems, which may overestimate the practical effectiveness of the vulnerabilities we identified. Regarding defensive approaches, our work lacks comprehensive solutions to the vulnerabilities demonstrated. While we discussed several potential defense strategies in the next section, we do not provide thoroughly tested, reliable defense methods with rigorous evaluation of their effectiveness or practical implementability. Furthermore, we have not explored the deeper theoretical principles underlying these vulnerabilities, limiting the ability to provide principled guidance for designing inherently robust depth estimation systems.
@@ Appendix E Broader Impact
We acknowledge there are legitimate concerns about how this research could be misused, thus we carefully considered the ethical implications of revealing vulnerabilities in stereo depth estimation systems. Malicious actors might exploit the vulnerabilities we have identified to compromise autonomous navigation systems in vehicles or robots, potentially causing accidents or enabling theft/tampering of autonomous systems. As a result, we implemented several safety controls to minimize misuse risk, including: ❶ limited disclosure of specific technical details that could enable immediate exploitation; ❷ establishment of a reasonable timeline for patches before full disclosure; ❸ creation of a centralized database of proposed attacks accessible only to verified researchers and industry partners We recommend similar safeguards for related research, including mandatory ethics review for attack demonstrations, implementation of differential privacy techniques to limit what information is shared, developing standardized responsible disclosure protocols specific to vision system vulnerabilities. By maintaining these ethical standards and safety controls, we can continue advancing security research while minimizing potential harm to systems that increasingly underpin critical infrastructure and everyday technologies.

### v2

@@ 6.5 Ablation Study
Figure 9: Attack performance of DepthVanish against RAFT-Stereo under different $\alpha$ and $\beta$ on KITTI dataset.
In this section, we conduct ablation analysis on the DepthVanish attack to assess the impact of the hyperparameters $\alpha$ and $\beta$ in the objective function of Eq. (9). As shown in Fig. 9, both parameters are critical for optimal attack performance. Specifically, it can be seen that the performance degraded significantly when $\alpha=0$, i.e., the $\mathcal{L}_{entropy}$ is removed from Eq. (9), which highlights the importance of the clear interval spacing for attack effectiveness. Moreover, the total variation constraint $\mathcal{L}_{tv}$ is also important where a clear performance degradation can be observed when $\beta$ decreases below 9. In summary, the synergistic combination of entropy and total variation regularization effectively ensures that our DepthVanish patches achieve the maximal attack performance
@@ 7 Conclusion
In this work, we present DepthVanish, a significant advancement in physical adversarial attack that jointly optimizes both texture element and interval structure of a patch to fool stereo depth estimation systems. By thoroughly analyzing the influence of regular spacing on naive texture repetition, we introduce a novel insight into enhancing the attack effectiveness and digital-to-physical transferability of the patch. To demonstrate the potentially dangerous consequences of depth estimation failure, we design the patch to be "disappear", where the patch is estimated as far away despite being physically close. Unlike previous methods limited to digital environments, our approach succeeds in both digital and physical settings, when evaluated against widely applied depth estimation models and commercial RGB-D cameras. These findings reveal critical vulnerabilities in current depth estimation technologies and raise concerns about their reliability in safety-critical autonomous systems.
Acknowledgments and Disclosure of Funding
This research was supported by Shenzhen Science and Technology Program (No. JCYJ20240813114237048), "Science and Technology Yongjiang 2035" key technology breakthrough plan project (No. 2025Z053). This research is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG4-GC-2023-008-1B), and National Research Foundation, Singapore and Infocomm Media Development Authority under its Trust Tech Funding Initiative. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore, Cyber Security Agency of Singapore, and Infocomm Media Development Authority. This work is also supported in part by Canada CIFAR AI Chairs Program, the Natural Sciences and Engineering Research Council of Canada, and JST-Mirai Program Grant No.JPMJMI20B8, JSPS KAKENHI Grant No.JP21H04877, No.JP23H03372, No.JP24K02920.
@@ Appendix A Experimental Environment
All of the experiments are conducted on a server with AMD EPYC 9554 64-core Processor and an NVIDIA L40 GPU, running Ubuntu 22.04. During the physical evaluation, the patch is printed out with an EPSON L18050 printer, i.e., a patch of physical size $(u,v)=(0.891m,1.26m)$ is printed out by filling A3 size paper.
@@ B.1 Calculation for Digital Deployment
During the digital experiments, the patches are first digitally deployed into the scene. As our aim is to achieve physical attack, the digital deployment is required to follow the physical constraints. Fortunately, such a physical-constrained digital patch deployment can be realized with the calibration information provided by the dataset. We show the detail calculation for the KITTI dataset below.
For KITTI dataset, we suppose the patch’s physical size and depth are predefined with regard to the camera 0 (the reference camera). Given the patch with physical size of $(w_{p},h_{p})$, we intend to deploy it into a KITTI scene with physical depth $e$. 1. Get the corresponding calibration information for the scene from ‘calib_cam_to_cam’ folder provided by KITTI scene flow dataset. 2. Retrieve the three rectified calibration matrix P_rect_02P\_rect\_02, P_rect_03P\_rect\_03, R_rect_00R\_rect\_00. 3. Specify the physical shifting $(x_{shift},y_{shift})$ (in meters) of the patch center with regard to the camera 0 principal axis. 4. Set the homogeneous coordinates for the corners of the patch as: $top\_left=(-w_{p}/2+x_{shift},-h_{p}/2+y_{shift},e,1)$, $top\_right=(w_{p}/2+x_{shift},-h_{p}/2+y_{shift},e,1)$, $bottom\_left=(-w_{p}/2+x_{shift},h_{p}/2+y_{shift},e,1)$, $bottom\_right=(-w_{p}/2+x_{shift},h_{p}/2+y_{shift},e,1)$. 5. For the pixel coordinates of the patch’s corner in the right stereo image, get them with $(P\_rect\_02\times R\_rect\_00)\cdot c$ where $c$ is the corners homogeneous coordinates. 6. Similarly, get the corresponding pixel coordinates for the left stereo image with $(P\_rect\_03\times R\_rect\_00)\cdot c$. 7. Finally, the patch is deployed into the scene by applying perspective transformation to fit the patch into the calculated pixel region.
For the calculation of DrivingStereo, the process is the same as KITTI except for the reference camera is camera 1 thus P_rect_101P\_rect\_101, P_rect_103P\_rect\_103 and R_rect_101R\_rect\_101 from their calibration file is adopted.
@@ B.2 Physical Stereo Image Capture
During the experiments of physical evaluation of our patch, we have to manually capture the scene due to the lack of RGB stereo capturing device. We adopt the i3DStreoid mobile application which is specifically designed to facilitate the capturing of stereo images with mobile phone of model iPhone 14 pro. In specific, a cutting board of A3 size is utilized to place the mobile phone for a $30cm$ baseline simulation. Then at each place a picture is taken by the mobile phone as one of the stereo image. However, we are aware of the inaccuracy of this stereo image capturing process, which is why we further test our patch with a commercial RGB-D camera, i.e., the IntelRealSense D435i.
@@ B.3 Setting for Attack Targets
For the PSMNet, DeepPruner and AANet, we follow the setup of Stereoscopic [1] and directly adopt the API provided at https://github.com/alexklwong/stereoscopic-universal-perturbations.git. As for the RAFT-Stereo and STTR, we use the official code release and integrate them following Stereoscopic code. And the checkpoint pretrianed on KITTI dataset for both models, i.e., ‘raftstereo-sceneflow.pth’ and ‘kitti_finetuned_model.pth.tar’, are adopt for experiments. Note that we aet $k=3$ for D1-error metric calculation during the evaluation for all the attack taregt models.
@@ C.1 Optimize under Different Setup
We conduct experiments to test the influence of the patch physical size and depth. As shown in Fig. 2, we test physical depth ranges from $5m$ to $21m$ and different physical size by setting the patch as a scaled size of $(0.891m,1.26m)$. In general, we observe that the patch remains effective across different physical sizes and depths. Empirically, a patch causes significant disturbance when the D1-error is larger than 80, which physically represents a disappeared region to the stereo system.
@@ C.2 Attack Transferability Results
As we have shown, our DepthVanish patch enjoys strong attack transferability over different stereo estimation models, we further visualize such merit in Fig. 3 and Fig. 4 on KITTI and DrivingStereo dataset respectively. Although there are some less effective results on the KITTI dataset, we observe that our DepthVanish patch successfully attacks all models, with varying degrees of transferability.
@@ Appendix D Limitation
Our research has several important limitations regarding the attack methodology. Due to a limited testing scope, we cannot guarantee that our attack methods are robust against the full spectrum of existing DNN architectures for stereo depth estimation. Additionally, we have not evaluated our attacks against existing robustified methods or defense mechanisms that may be implemented in real-world systems, which may overestimate the practical effectiveness of the vulnerabilities we identified. Regarding defensive approaches, our work lacks comprehensive solutions to the vulnerabilities demonstrated. While we discussed several potential defense strategies in the next section, we do not provide thoroughly tested, reliable defense methods with rigorous evaluation of their effectiveness or practical implementability. Furthermore, we have not explored the deeper theoretical principles underlying these vulnerabilities, which limits our ability to provide principled guidance for designing inherently robust depth estimation systems.
@@ Appendix E Broader Impact
We acknowledge there are legitimate concerns about how this research could be misused, thus we carefully considered the ethical implications of revealing vulnerabilities in stereo depth estimation systems. Malicious actors might exploit the vulnerabilities we have identified to compromise autonomous navigation systems in vehicles or robots, potentially causing accidents or enabling theft/tampering of autonomous systems. As a result, we implemented several safety controls to minimize misuse risk, including: ❶ limited disclosure of specific technical details that could enable immediate exploitation; ❷ establishment of a reasonable timeline for patches before full disclosure; ❸ creation of a centralized database of proposed attacks accessible only to verified researchers and industry partners We recommend similar safeguards for related research, including mandatory ethics review for attack demonstrations, the implementation of differential privacy techniques to limit what information is shared, and the development of standardized responsible disclosure protocols specific to vision system vulnerabilities. By maintaining these ethical standards and safety controls, we can continue advancing security research while minimizing potential harm to systems that increasingly underpin critical infrastructure and everyday technologies.